OpenStudy (anonymous):

round 8 5/9 to the nearest whole number. 7 8 9 10

OpenStudy (michele_laino):

please what is \[8 + \frac{5}{9}\] ?

OpenStudy (michele_laino):

hint: \[8 + \frac{5}{9} = \frac{{\left( {8 \cdot 9} \right) + 5}}{9}\]